# Sylvester domains and pro-$p$ groups

### v1

@@ 7. The equality $\operatorname{rk}_{G}=\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}$
Since $l_{G}(M)$ is finite by Claim 7.4, we have $\lim_{i\to\infty}\frac{l_{G}(M)}{i}=0$, which concludes the proof. ∎
Now we are ready to finish the proof of Theorem 1.1.
Proof of Theorem 1.1.
By Lemma 4.6 and Corollary 3.8, we can assume that $G$ is a mild flag free-by-$\mathbb{Z}_{p}$ pro-$p$ group. Let $\mathcal{Q}$ and $\mathbb{D}=\mathcal{D}[\![t]\!][\![s;\sigma,\delta]\!]$ be as in Section 5. Observe that the valuation $w$ on $\mathbb{D}$ naturally extends to $\mathcal{Q}$ and the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{Q}$ is continuous by Lemma 5.1. By Theorem 7.1, $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}\leq\operatorname{rk}_{G}\,$. In the proof of Theorem 6.13 we showed that $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}=\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}$. Since $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}\geq\operatorname{rk}_{G}$, we have that $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}=\operatorname{rk}_{G}$. ∎
@@ 8. Abstract subgroups of free-by-$\mathbb{Z}_{p}$ pro-$p$ groups
If $K$ is of characteristic 0, the parts (1) and (2) of Conjecture 2 are known to be true [JZ19a], and for part (3) we know that there exists an embedding $\varphi:K[\Gamma]\to\mathcal{Q}$ into a division ring $\mathcal{Q}$ such that $\lim_{i\to\infty}\operatorname{rk}_{\Gamma/N_{i}}(A)=\operatorname{rk}_{\mathcal{Q}}(\varphi(A))$ [JZLA20], but we still do not know whether $\varphi$ is universal in general [JZ21]. If $K$ is of positive characteristic, the parts (1), (2) and (3) are only known when $\Gamma$ is amenable ([JZ19b, JZ21]).
Let now $G$ be a free-by-$\mathbb{Z}_{p}$ pro-$p$ group and $\Gamma$ a finitely generated (abstract) subgroup of $G$. It is clear that $\Gamma$ is locally indicable and from [JZ21, Theorem 3.7] we know that there exist a universal division $\mathbb{F}_{p}[\Gamma]$-ring of fractions $\mathcal{D}_{\mathbb{F}_{p}[\Gamma]}$. The following theorem provides a particular case of Conjecture 2 for $\Gamma$.
@@
Claim 8.2.
$\mathcal{D}_{H}$ is the universal division ring of fraction of $\mathbb{F}_{p}[H]$
Proof.
This is proved in [JZ24, Proposition 3.5]. ∎
Claim 8.3.
The induced map $\mathcal{D}_{H}*(\Gamma/H)\to R$ is an isomorphism.
Proof.
Let $\gamma_{1},\ldots,\gamma_{n}$ lie in different clases modulo $H$. There exists an open normal subgroup $U$ of $G$ containing $N$ such that $\gamma_{1},\ldots,\gamma_{n}$ lie in different clases modulo $U$. By Proposition 3.4(d), $\sum_{i=1}^{n}\mathcal{D}_{H}\gamma_{i}$ is direct. This implies the claim. ∎
The two claims toghether imply that the ring $R$ is isomorphic to a crossed product $\mathcal{D}_{\mathbb{F}_{p}[H]}*(\Gamma/H)$. Since $\Gamma/H$ is finitely generated, $R$ a Noetherian ring. Thus, $\mathcal{D}_{\Gamma}$ is isomorphic to the Ore division ring of fractions of $R$. Hence, we obtain that $\mathcal{D}_{\Gamma}$ will be the universal division ring of fraction of $\mathbb{F}_{p}[\Gamma]$. ∎
@@ Appendix A Localization of profinite rings
J. Beachy has shown in [Bea93] that $R_{\Sigma}$ and $\lambda$ can be constructed as follows. From now on we assume that $\Sigma$ is complete. Let $R^{n}$ denote the rows of $n$ elements in $R$, ${}^{n}R$ the columns of $n$ elements in $R$ and $\Sigma_{n}$ denote the subset of $n$-by-$n$ matrices in $\Sigma$. Consider the set $T_{\Sigma}$ given by the disjoint union of all the products $R^{n}\times\Sigma_{n}\times{{}^{n}R}$, that is, the triples $(a,C,x)$ where $a$ is a row matrix over $R$ of length $n$, $C\in\Sigma_{n}$ and $x$ is a column matrix over $R$ of height $n$. We define two binary operations of sum and product in $T_{\Sigma}$ as follows:
@@
Let $\Sigma^{-1}_{0}R$ be the subsemigroup of $\Sigma^{-1}R$ generated by all the elements of the form $(a\colon C\colon 0)$ and $(0\colon C\colon y)$. It is shown in [Bea93] that all the elements of $\Sigma^{-1}_{0}R$ are either of that form or of the form $\left(\begin{pmatrix}a&0\end{pmatrix}\colon\begin{pmatrix}A&0\\ 0&D\end{pmatrix}\colon\begin{pmatrix}0\\ x\end{pmatrix}\right)$ for blocks of the appropriate sizes. This subsemigroup allows us to define the congruence relation $\mathcal{R}_{2}\subseteq\Sigma^{-1}R\times\Sigma^{-1}R$ on $\Sigma^{-1}R$, where two elements $\overline{p}\,,\overline{q}\in\Sigma^{-1}R$ are equivalent if there exists $\overline{z}_{1}\,,\overline{z}_{2}\in\Sigma_{0}^{-1}R$ such that $\overline{p}+\overline{z}_{1}=\overline{q}+\overline{z}_{2}$.
The quotient space $\Sigma^{-1}R/\sim_{\mathcal{R}_{2}}$ inherits the sum and product, under which it becomes an associative ring with trivial element being the class [(1:1:0)][(1:1:0)][ ( 1 : 1 : 0 ) ] and the identity element being the class [(1:1:1)][(1:1:1)][ ( 1 : 1 : 1 ) ]. The additive inverse of $[(a\colon C\colon x)]$ is $[(a\colon C\colon-x)]$. One of the main results of [Bea93] is that this quotient space is precisely the universal localization $R_{\Sigma}$, where the map $\lambda\colon R\to R_{\Sigma}$ is given by $x\mapsto[(1\colon 1\colon x)]$.
@@
[…]

### v2

@@ 7. The equality $\operatorname{rk}_{G}=\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}$
Since $l_{G}(M)$ is finite by Lemma 7.4, we have $\lim_{i\to\infty}\frac{l_{G}(M)}{i}=0$, which concludes the proof.
Now we are ready to finish the proof of Theorem 1.1. By Lemma 4.6 and Corollary 3.6, we can assume that $G$ is a mild flag free-by-$\mathbb{Z}_{p}$ pro-$p$ group. Let $\mathcal{Q}$ and $\mathbb{D}=\mathcal{D}[\![t]\!][\![s;\sigma,\delta]\!]$ be as in Section 5. Observe that the valuation $w$ on $\mathbb{D}$ naturally extends to $\mathcal{Q}$ and the embedding $\mathbb{F}_{p}[\![G]\!]\to\mathcal{Q}$ is continuous by Lemma 5.1. By Theorem 7.1, $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}\leq\operatorname{rk}_{G}\,$. In the proof of Theorem 6.13 we showed that $\varphi^{\#}\operatorname{rk}_{\mathcal{Q}}=\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}$. Since $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}\geq\operatorname{rk}_{G}$, we have that $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}=\operatorname{rk}_{G}$. ∎
@@ 8. Abstract subgroups of free-by-$\mathbb{Z}_{p}$ pro-$p$ groups
If $K$ is of characteristic 0, the parts (1) and (2) of Conjecture 2 are known to be true [23], and for part (3) we know that there exists an embedding $\varphi:K[\Gamma]\to\mathcal{Q}$ into a division ring $\mathcal{Q}$ such that $\lim_{i\to\infty}\operatorname{rk}_{\Gamma/N_{i}}(A)=\operatorname{rk}_{\mathcal{Q}}(\varphi(A))$ [19], but we still do not know whether $\varphi$ is universal in general [25]. If $K$ is of positive characteristic, the parts (1), (2) and (3) are only known when $\Gamma$ is amenable ([22, 25]).
Let now $G$ be a free-by-$\mathbb{Z}_{p}$ pro-$p$ group and $\Gamma$ a finitely generated (abstract) subgroup of $G$. It is clear that $\Gamma$ is locally indicable and from [25, Theorem 3.7] we know that there exist a universal division $\mathbb{F}_{p}[\Gamma]$-ring of fractions $\mathcal{D}_{\mathbb{F}_{p}[\Gamma]}$. The following theorem provides a particular case of Conjecture 2 for $\Gamma$.
@@
In [26, Prop. 3.5], it was shown that $\mathcal{D}_{H}$ is the universal division ring of fractions of $\mathbb{F}_{p}[H]$. We also claim that the induced surjective map $\mathcal{D}_{H}*(\Gamma/H)\to R$ is an isomorphism. Let $\gamma_{1},\ldots,\gamma_{n}$ lie in different classes modulo $H$. There exists an open normal subgroup $U$ of $G$ containing $N$ such that $\gamma_{1},\ldots,\gamma_{n}$ lie in different clases modulo $U$. By Proposition 3.4(d), $\sum_{i=1}^{n}\mathcal{D}_{H}\gamma_{i}$ is direct. Hence, it sends linearly independent elements to linearly independent elements, establishing its injectivity.
These two facts together imply that the ring $R$ is isomorphic to a crossed product $\mathcal{D}_{\mathbb{F}_{p}[H]}*(\Gamma/H)$. Since $\Gamma/H$ is finitely generated, $R$ a Noetherian ring. Thus, $\mathcal{D}_{\Gamma}$ is isomorphic to the Ore division ring of fractions of $R$. Hence, we obtain that $\mathcal{D}_{\Gamma}$ will be the universal division ring of fraction of $\mathbb{F}_{p}[\Gamma]$. ∎
@@ Appendix A Localization of profinite rings
J. Beachy has shown in [6] that $R_{\Sigma}$ and $\lambda$ can be constructed as follows. From now on we assume that $\Sigma$ is complete. Let $R^{n}$ denote the rows of $n$ elements in $R$, $^{n}R$ the columns of $n$ elements in $R$ and $\Sigma_{n}$ denote the subset of $n$-by-$n$ matrices in $\Sigma$. Consider the set $T_{\Sigma}$ given by the disjoint union of all the products $R^{n}\times\Sigma_{n}\times{{}^{n}R}$, that is, the triples $(a,C,x)$ where $a$ is a row matrix over $R$ of length $n$, $C\in\Sigma_{n}$ and $x$ is a column matrix over $R$ of height $n$. We define two binary operations of sum and product in $T_{\Sigma}$ as follows:
@@
Let $\Sigma^{-1}_{0}R$ be the subsemigroup of $\Sigma^{-1}R$ generated by all the elements of the form $(a\colon C\colon 0)$ and $(0\colon C\colon y)$. It is shown in [6] that all the elements of $\Sigma^{-1}_{0}R$ are either of that form or of the form $\left(\begin{pmatrix}a&0\end{pmatrix}\colon\begin{pmatrix}A&0\\ 0&D\end{pmatrix}\colon\begin{pmatrix}0\\ x\end{pmatrix}\right)$ for blocks of the appropriate sizes. This subsemigroup allows us to define the congruence relation $\mathcal{R}_{2}\subseteq\Sigma^{-1}R\times\Sigma^{-1}R$ on $\Sigma^{-1}R$, where two elements $\overline{p}\,,\overline{q}\in\Sigma^{-1}R$ are equivalent if there exists $\overline{z}_{1}\,,\overline{z}_{2}\in\Sigma_{0}^{-1}R$ such that $\overline{p}+\overline{z}_{1}=\overline{q}+\overline{z}_{2}$.
The quotient space $\Sigma^{-1}R/\sim_{\mathcal{R}_{2}}$ inherits the sum and product, under which it becomes an associative ring with trivial element being the class $[(1:1:0)]$ and the identity element being the class $[(1:1:1)]$. The additive inverse of $(a\colon C\colon x)]$ is $(a\colon C\colon-x)]$. One of the main results of [6] is that this quotient space is precisely the universal localization $R_{\Sigma}$, where the map $\lambda\colon R\to R_{\Sigma}$ is given by $x\mapsto[(1\colon 1\colon x)]$.
@@
We have that $\{[(1\colon 1\colon 0)]\}$ is closed in $R_{\Sigma}$.